# Impact of memory on clustering in spontaneous particle aggregation

### v2

@@ 1 Introduction
Models with memory and delays are ubiquitous in explaining collective behaviour of organisms, ranging in scale from bacterial and amoeboid chemotaxis [Berg:1972:CEC, Erban:2007:TEA] to behaviour of social insects and flocks of birds [Sumpter:2006:PCA, Sumpter:2010:CAB]. In the case of bacteria, the memory is incorporated in purely biochemical terms, in the form of the signal transduction network capable of excitation and adaptation dynamics [Spiro:1997:MEA, Alon:1999:RBC]. This enables a cell to ‘memorize’ past environmental signals and compare to their current state to inform decision making [Berg:1990:BM]. In individual-based models, such memory effects can be accounted for by ordinary differential equation (ODE) or stochastic differential equation (SDE) models describing internal variables of each individual [Erban:2004:ICB, Erban:2007:TEA]. A typical example of such internal variables are concentrations of key intracellular biochemical species. At the collective (macroscopic) level, populations of bacteria or other cells can be described by partial differential equations (PDEs) describing the density of cells and key environmental signals [Erban:2005:STS, Erban:2020:SMR], where memory can also be interpreted as delays in the processing of some signals. Incorporating such delays into mathematical models of collective phenomena can provide more accurate explanations of behavioural properties of groups of animals [Sun:2014:TDC, Erban:2016:CSM], robots [Taylor:2015:MMT], or spreading processes in social and biological systems [Rosvall:2014]. Moreover, non-Markovian characteristics resulting from ordering of interactions in temporal complex networks were identified as an important mechanism that alters causality and affects dynamical processes in social, technological and biological systems [Scholtes2014, Williams2022]. Considering biological agents (cells or animals), a trade-off between having no memory at all and remembering too much (including very distant and irrelevant past states) has been achieved by evolution. For example, bacterial ‘biochemical memory’ of past environmental signals evolved to last for the duration of a few seconds [Berg:1990:BM].
@@
In this paper we focus on a model of spontaneous aggregation of animals or cells resulting from random diffusive motion of individuals [BHW:2012:PhysD, HO:2015:DCDSB]. The individuals respond to the local population density observed in their neighbourhood by increasing or decreasing the amplitude of their random motion. This kind of behaviour has been observed in insects, for instance, the pre-social German cockroach (B. germanica), which are known to be attracted to dark, warm and humid places. However, it has been shown that cockroach larvae also aggregate spontaneously [Jeanson:2005:AnimBehav], i.e., in the absence of any environmental template or heterogeneity. Moreover, this type of dynamics can also be used to describe formation of P Granules in C. elegans embryos resulting in spontaneous protein aggregation [Brangwynne:2009:Science]. We note that spontaneous aggregates form at random locations in space. If such spontaneous aggregates are to form on a regular lattice, additional interactions between individuals must be incorporated into the model [Lopez:2025:SCR].
@@
To incorporate memory into the first-order model of spontaneous aggregation [BHW:2012:PhysD], we introduce a set of $K\geq 1$ internal variables of each agent. Such internal variables can have different physical meanings in different biological models. In applications to bacterial chemotaxis, individuals compare current and past concentrations of extracellular signals, and internal variables can be identified with the concentration of molecules transducing the signal in detailed models of the chemotactic network [Celani:2010]. The internal variables can also be viewed as effectively storing the memory of the state of the external environment in the vicinity of the individual. In our case, we consider a model of the spontaneous aggregation (i.e., with no external chemical signal) and $K\geq 1$ internal variables store a time-dependent memory of how crowded the neighborhood was in the past. In particular, each internal variable describes a ‘layer’ of memory, which is subject to two effects: (i) production (or excitation) from the internal variable of the next higher order, and (ii) spontaneous decay with a constant rate. The internal variable of the highest (i.e., $K$-th) order is then subject to random excitation, with amplitude modulated by a nonlinear function of the number density of agents observed in the physical neighbourhood. We note that in the first-order model introduced in [BHW:2012:PhysD], the positions of the agents in the physical space were directly subjected to random excitation (Brownian motion). In particular, our investigation extends their model by introducing a chain of $K$ internal variables that allow the agents to ‘remember’ the densities they encountered in the past.
@@
Our results show that the introduction of memory with a few layers, $K$, leads to the formation of a smaller number of larger clusters of agents. This trend is observed until $K=3$ or $K=4$, depending on the spatial dimension of the studied system, as shown in Sections 6 and 7. When the number of memory layers $K$ is increased further, i.e., as the memory becomes ‘longer’, its effect starts to be disruptive. This is manifested by the increasing proportion of ‘outliers’, which are the agents that are not part of any cluster. We therefore conclude that short-term or medium-term memory enhances spontaneous aggregation, while long-term memory disrupts it. This observation is ecologically relevant because many organisms - from bacteria to insects to higher animals - do not respond to their environment instantaneously but instead use some form of memory or internal state to modulate their behavior. Species that benefit from aggregation (e.g., for warmth, protection, or reproduction) may have developed memory systems that retain only short-to-medium-term environmental information. Conversely, overly long memory may reduce responsiveness to current cues and hinder effective social or spatial organization. Therefore, organisms must balance between being too reactive (leading to unstable or erratic behavior) and being too inertial (failing to adapt to rapidly changing environments). This is consistent with ecological and evolutionary studies that view memory as an adaptive trait shaped by natural selection, see, e.g., [Couzin:2002, Chen:2024].
@@
There is a vast body of literature focusing on mathematical models of aggregation dynamics and their analysis, with the number of papers counting in hundreds. These models include not only aggregation of agents or particles in the physical space (e.g., [Jeanson:2005:AnimBehav, Brangwynne:2009:Science, Lopez:2025:SCR, Keller:1970:JTB, Haskovec:2009:JSP]), but also in abstract phase space, i.e., models of opinion formation [Hegselmann:2002:JASSS, Jabin:2014:JDE], flocking and swarming [Vicsek:1995:PRL, Cucker:2007:IEEE, Olfati:2006:IEEE] or pattern formation resulting from attraction-repulsion interactions [Dorsogna:2006:PRL, Carrillo:2019:AP]. For a relatively recent survey of models of collective motion we refer to [Vicsek:2012:PR]. Many variants of these models have been studied, including models with nonlocal interactions [Mogilner:1999:JMB, Bernoff:2013:SIAMRev, Carrillo:2015:KRM], topological interactions [Haskovec:2013:PhysD], hydrodynamic models [Choi:2019:SIMA], models with hierarchy [Shen:2007:SIAM, Voggt:2019:PhysRevRes] and leaders [Cicolani:2024:arXiv], models with noise [Cattiaux:2018:AAP] and delay [Choi:2021:MMAS, Haskovec:2021:SIADS] and combination of those [Erban:2016:CSM]. In [Ahn:2023:MBE] the authors considered a variant of the Cucker-Smale flocking model with a Caputo fractional derivative in the time variable. The presence of the fractional derivative can be interpreted as a form of memory. The authors derived sufficient conditions for asymptotic flocking, which occurs at an algebraic rate, in contrast to the exponential flocking asymptotics in the original Cucker-Smale model [Cucker:2007:IEEE]. Memory induced aggregation in collective foraging was studied in [Nauta:2020:HFP]. Using an agent-based model, the authors argued that aggregation around salient patches can occur through formation of collective memory realized through local interactions and global displacement using Lévy walks. The impact of emergent memory on opinion dynamics was analyzed in [Boschi:2021:PhysA], with the aim to explain how the exposure of the society to certain events deeply changes people’s perception of the present and future. An analytical way was proposed to measure how much information a society can remember when an extensive number of news items was presented.
@@
The paper is organized as follows. In Section 2 we describe the (first-order) spontaneous aggregation model and summarize its main properties studied in the literature [BHW:2012:PhysD]. In Section 3 we introduce memory into this model and infer the main mathematical properties of the model with memory. In Section 4 we derive the formal macroscopic limit of the system as the number of agents tends to infinity, obtaining the corresponding Fokker-Planck equation. In Section 5 we then characterize its steady states to gain an insight into the patterns (clusters) formed by the system. In Section 6 we report the results of the stochastic simulations of the individual-based model in the spatially one-dimensional setting, while the two-dimensional results are presented in Section 7. Here we also provide a statistical evidence that the long-term memory inhibits the particles’ responsiveness to environmental stimuli. We conclude with the discussion in Section 8.
@@ 2 Spontaneous aggregation model without memory
The individual-based stochastic model introduced in reference [BHW:2012:PhysD] under the name ‘direct aggregation model’ consists of a group of $N\geq 2$ biological agents (cells or animals), characterized by their positions ${\mathbf{x}}_{i}(t)\in\mathbb{R}^{d}$, with spatial dimension $d\in\{1,2,3\}$ and $i\in[N]$, where we have denoted the set of indices by $N]:=\{1,2,\ldots,N\}$. Every individual senses the average density of its close neighbours, given by
@@
In reference [BHW:2012:PhysD], the mean-field limit (as $N\to\infty$) of the individual-based model (2.1)–(2.3) is studied, describing the system in terms of the number particle density $\varrho\equiv\varrho(t,{\mathbf{x}})$ with ${\mathbf{x}}\in\mathbb{R}^{d}$. The time evolution of the density $\varrho$ is subject to
@@
Here we identify the convection term $\nabla\cdot\left(\varrho\,\nabla\big[G(W\ast\varrho)^{2}\big]\right)$ as the ‘driving force’ responsible for the eventual formation of aggregates (recall the monotonicity assumption on $G$). Then the following condition for aggregation, emanating from a perturbation of a given constant steady state $\varrho\equiv\varrho_{0}>0$, can be derived [BHW:2012:PhysD],
@@ 3 Spontaneous aggregation model with memory
The model with memory (3.1) has been introduced in a way that it can be reduced to the original spontaneous aggregation model (2.1)–(2.3) if we set $\alpha_{k}:=1$ for all $k\in[K]$ and pass to the limit $\varepsilon_{k}\to 0$ for all $k\in[K]$. However, it can be also formulated in a more general way, which provides connections with the memory models in the literature [Fagan:2023:RDM, Nauta:2020:HFP, Kim:2024:ISM]. Denoting the $j$-th component of ${\mathbf{y}}_{i}^{k}\in\mathbb{R}^{d}$ by $y_{i,j}^{k}$, for $i\in[N]$, $j\in[d]$ and $k\in[K]$, we introduce the following notation
@@
The reformulation of model (3.1) into the SDE form (3.4) aligns it more closely with the formulation of the reinforced diffusion model for memory-mediated animal movement [Fagan:2023:RDM]. The key difference in our study is that we focus on spontaneous aggregation, whereas some models in the literature examine aggregation in response to environmental cues to locate specific targets [Nauta:2020:HFP, Kim:2024:ISM]. Additionally, we explicitly describe the memory variables, whereas some models introduce memory implicitly through delays [Fagan:2023:RDM]. Model (3.4) can also be rewritten in the delay form, because the solution of the SDE system (3.4) is given by the variation-of-constants formula [Mao:2007, Section 3.3]
@@
where ${\mathbf{y}}_{i,j}(0)$ is a prescribed initial condition at time $t=0$. The SDE form of the memory model (3.4) can, in general, be used to describe the ‘chemical memory’ of unicellular organisms, where internal variables ${\mathbf{y}}_{i}^{k}$ can be identified with concentrations of key chemical species of signal transduction networks [Erban:2005:STS, Erban:2007:TEA]. The behaviour of such memory models will depend on the properties of matrix $A^{(K)}\in\mathbb{R}^{K\times K}$, on its eigenvalues and eigenvectors. In this paper, we use the specific form of the general model (3.4) given by (3.5). In particular, choosing ${\mathbf{y}}_{i,j}(0)=0$ in equation (3.6), we have for $y_{i,j}^{1}\equiv y_{i,j}^{1}(t)$
@@
To understand how the choice of the number of layers $K$ impacts the ‘length’ of the memory, we calculate the mean of the normalized version of the kernel $\kappa=\kappa(t)$,
@@
We observe that the length of the memory increases linearly with $K$. In Sections 6 and 7, we will numerically investigate the impact of the choice of $K$ on the clustering properties of the individual-based model with memory given by equations (2.1) and (3.9). Let us note that the kernel (3.8) takes very different shapes even if the length (3.10) is kept constant. For instance, in Figure 1 we plot $\kappa=\kappa(t)$ for the length
@@
Finally, let us note that this individual-based model can be also viewed for $K=1$ as a simplified version of what is called ‘the second-order model’ in reference [BHW:2012:PhysD]. Indeed, if $K=1$, then the SDE system (3.1) reduces to
@@ 4 Formal derivation of the Fokker-Planck equation
The formal large population limit as $N\to\infty$ of the system (3.1) with the average density given by (2.1) can be derived by a generalization of the procedure used in [BHW:2012:PhysD, Section 3]. Since some nontrivial steps were omitted in [BHW:2012:PhysD], we include the detailed formal derivation here. The main idea is to ‘linearize’ (3.1) by introducing the average densities $\vartheta_{i}=\vartheta_{i}(t)$ for $i\in[N]$ as a set of $N$ independent variables. According to (2.1) and the first equation of (3.1), these new variables are subject to the system of differential equations
@@
One then applies the Itô formula [Mao:2007] to write down the Liouville equation [Cercignani:1988] for the $N$-particle density $f^{N}=f^{N}(t,{\mathbf{x}}_{1},\dots,{\mathbf{x}}_{N},{\mathbf{y}}^{1}_{1},\dots,{\mathbf{y}}^{1}_{N},\dots,{\mathbf{y}}^{K}_{1},\dots,{\mathbf{y}}^{K}_{N},\vartheta_{1},\dots,\vartheta_{N})$,

### v3

@@ 1 Introduction
Models with memory and delays are ubiquitous in explaining collective behaviour of organisms, ranging in scale from bacterial and amoeboid chemotaxis [3, 26] to behaviour of social insects and flocks of birds [53, 54]. In the case of bacteria, the memory is incorporated in purely biochemical terms, in the form of the signal transduction network capable of excitation and adaptation dynamics [52, 2]. This enables a cell to ‘memorize’ past environmental signals and compare to their current state to inform decision making [4]. In individual-based models, such memory effects can be accounted for by ordinary differential equation (ODE) or stochastic differential equation (SDE) models describing internal variables of each individual [24, 26]. A typical example of such internal variables are concentrations of key intracellular biochemical species. At the collective (macroscopic) level, populations of bacteria or other cells can be described by partial differential equations (PDEs) describing the density of cells and key environmental signals [25, 21], where memory can also be interpreted as delays in the processing of some signals. Incorporating such delays into mathematical models of collective phenomena can provide more accurate explanations of behavioural properties of groups of animals [55, 22], robots [56], or spreading processes in social and biological systems [49]. Moreover, non-Markovian characteristics resulting from ordering of interactions in temporal complex networks were identified as an important mechanism that alters causality and affects dynamical processes in social, technological and biological systems [50, 60]. Considering biological agents (cells or animals), a trade-off between having no memory at all and remembering too much (including very distant and irrelevant past states) has been achieved by evolution. For example, bacterial ‘biochemical memory’ of past environmental signals evolved to last for the duration of a few seconds [4].
@@
In this paper we focus on a model of spontaneous aggregation of animals or cells resulting from random diffusive motion of individuals [8, 34]. The individuals respond to the local population density observed in their neighbourhood by increasing or decreasing the amplitude of their random motion. This kind of behaviour has been observed in insects, for instance, the pre-social German cockroach (B. germanica), which are known to be attracted to dark, warm and humid places. However, it has been shown that cockroach larvae also aggregate spontaneously [40], i.e., in the absence of any environmental template or heterogeneity. Moreover, this type of dynamics can also be used to describe formation of P Granules in C. elegans embryos resulting in spontaneous protein aggregation [7]. We note that spontaneous aggregates form at random locations in space. If such spontaneous aggregates are to form on a regular lattice, additional interactions between individuals must be incorporated into the model [44].
@@
To incorporate memory into the first-order model of spontaneous aggregation [8], we introduce a set of $K\geq 1$ internal variables of each agent. Such internal variables can have different physical meanings in different biological models. In applications to bacterial chemotaxis, individuals compare current and past concentrations of extracellular signals, and internal variables can be identified with the concentration of molecules transducing the signal in detailed models of the chemotactic network [12]. The internal variables can also be viewed as effectively storing the memory of the state of the external environment in the vicinity of the individual. In our case, we consider a model of the spontaneous aggregation (i.e., with no external chemical signal) and $K\geq 1$ internal variables store a time-dependent memory of how crowded the neighborhood was in the past. In particular, each internal variable describes a ‘layer’ of memory, which is subject to two effects: (i) production (or excitation) from the internal variable of the next higher order, and (ii) spontaneous decay with a constant rate. The internal variable of the highest (i.e., $K$-th) order is then subject to random excitation, with amplitude modulated by a nonlinear function of the number density of agents observed in the physical neighbourhood. We note that in the first-order model introduced in [8], the positions of the agents in the physical space were directly subjected to random excitation (Brownian motion). In particular, our investigation extends their model by introducing a chain of $K$ internal variables that allow the agents to ‘remember’ the densities they encountered in the past.
@@
Our results show that the introduction of memory with a few layers, $K$, leads to the formation of a smaller number of larger clusters of agents. This trend is observed until $K=3$ or $K=4$, depending on the spatial dimension of the studied system, as shown in Sections 6 and 7. When the number of memory layers $K$ is increased further, i.e., as the memory becomes ‘longer’, its effect starts to be disruptive. This is manifested by the increasing proportion of ‘outliers’, which are the agents that are not part of any cluster. We therefore conclude that short-term or medium-term memory enhances spontaneous aggregation, while long-term memory disrupts it. This observation is ecologically relevant because many organisms - from bacteria to insects to higher animals - do not respond to their environment instantaneously but instead use some form of memory or internal state to modulate their behavior. Species that benefit from aggregation (e.g., for warmth, protection, or reproduction) may have developed memory systems that retain only short-to-medium-term environmental information. Conversely, overly long memory may reduce responsiveness to current cues and hinder effective social or spatial organization. Therefore, organisms must balance between being too reactive (leading to unstable or erratic behavior) and being too inertial (failing to adapt to rapidly changing environments). This is consistent with ecological and evolutionary studies that view memory as an adaptive trait shaped by natural selection, see, e.g., [18, 14].
@@
There is a vast body of literature focusing on mathematical models of aggregation dynamics and their analysis, with the number of papers counting in hundreds. These models include not only aggregation of agents or particles in the physical space (e.g., [40, 7, 44, 41, 35]), but also in abstract phase space, i.e., models of opinion formation [38, 39], flocking and swarming [57, 19, 48] or pattern formation resulting from attraction-repulsion interactions [20, 9]. For a relatively recent survey of models of collective motion we refer to [58]. Many variants of these models have been studied, including models with nonlocal interactions [46, 5, 10], topological interactions [36], hydrodynamic models [15], models with hierarchy [51, 59] and leaders [17], models with noise [11] and delay [16, 37] and combination of those [22]. In [1] the authors considered a variant of the Cucker-Smale flocking model with a Caputo fractional derivative in the time variable. The presence of the fractional derivative can be interpreted as a form of memory. The authors derived sufficient conditions for asymptotic flocking, which occurs at an algebraic rate, in contrast to the exponential flocking asymptotics in the original Cucker-Smale model [19]. Memory induced aggregation in collective foraging was studied in [47]. Using an agent-based model, the authors argued that aggregation around salient patches can occur through formation of collective memory realized through local interactions and global displacement using Lévy walks. The impact of emergent memory on opinion dynamics was analyzed in [6], with the aim to explain how the exposure of the society to certain events deeply changes people’s perception of the present and future. An analytical way was proposed to measure how much information a society can remember when an extensive number of news items was presented.
@@
The paper is organized as follows. In Section 2 we describe the (first-order) spontaneous aggregation model and summarize its main properties studied in the literature [8]. In Section 3 we introduce memory into this model and infer the main mathematical properties of the model with memory. In Section 4 we derive the formal macroscopic limit of the system as the number of agents tends to infinity, obtaining the corresponding Fokker-Planck equation. In Section 5 we then characterize its steady states to gain an insight into the patterns (clusters) formed by the system. In Section 6 we report the results of the stochastic simulations of the individual-based model in the spatially one-dimensional setting, while the two-dimensional results are presented in Section 7. Here we also provide a statistical evidence that the long-term memory inhibits the particles’ responsiveness to environmental stimuli. We conclude with the discussion in Section 8.
@@ 2 Spontaneous aggregation model without memory
The individual-based stochastic model introduced in reference [8] under the name ‘direct aggregation model’ consists of a group of $N\geq 2$ biological agents (cells or animals), characterized by their positions ${\mathbf{x}}_{i}(t)\in\mathbb{R}^{d}$, with spatial dimension $d\in\{1,2,3\}$ and $i\in[N]$, where we have denoted the set of indices by $N]:=\{1,2,\ldots,N\}$. Every individual senses the average density of its close neighbours, given by
@@
In reference [8], the mean-field limit (as $N\to\infty$) of the individual-based model (2.1)–(2.3) is studied, describing the system in terms of the number particle density $\varrho\equiv\varrho(t,{\mathbf{x}})$ with ${\mathbf{x}}\in\mathbb{R}^{d}$. The time evolution of the density $\varrho$ is subject to
@@
Here we identify the convection term $\nabla\cdot\left(\varrho\,\nabla\big[G(W\ast\varrho)^{2}\big]\right)$ as the ‘driving force’ responsible for the eventual formation of aggregates (recall the monotonicity assumption on $G$). Then the following condition for aggregation, emanating from a perturbation of a given constant steady state $\varrho\equiv\varrho_{0}>0$, can be derived [8],
@@ 3 Spontaneous aggregation model with memory
The internal variables ${\mathbf{y}}_{i}^{k}={\mathbf{y}}_{i}^{k}(t)\in\mathbb{R}^{d}$ represent successive layers of biological memory through which information about past local crowding is processed before affecting movement. Rather than storing the past density explicitly, the agent integrates past sensory input through a cascade of internal states with finite relaxation times. Such cascades naturally arise in biological systems, for instance in biochemical signalling networks where an external stimulus is transmitted through multiple molecular intermediates, or in neural or behavioral processing where information is gradually filtered and attenuated.
In this interpretation, the parameter $K$ does not represent distinct memories of specific past times, but rather the depth of the memory-processing cascade. Increasing $K$ increases the number of intermediate steps through which information must pass before influencing motion, thereby extending the effective memory timescale and smoothing the response to rapid environmental fluctuations.
The model with memory (3.1) has been introduced in a way that it can be reduced to the original spontaneous aggregation model (2.1)–(2.3) if we set $\alpha_{k}:=1$ for all $k\in[K]$ and pass to the limit $\varepsilon_{k}\to 0$ for all $k\in[K]$. However, it can be also formulated in a more general way, which provides connections with the memory models in the literature [31, 47, 42]. Denoting the $j$-th component of ${\mathbf{y}}_{i}^{k}\in\mathbb{R}^{d}$ by $y_{i,j}^{k}$, for $i\in[N]$, $j\in[d]$ and $k\in[K]$, we introduce the following notation
@@
The reformulation of model (3.1) into the SDE form (3.4) aligns it more closely with the formulation of the reinforced diffusion model for memory-mediated animal movement [31]. The key difference in our study is that we focus on spontaneous aggregation, whereas some models in the literature examine aggregation in response to environmental cues to locate specific targets [47, 42]. Additionally, we explicitly describe the memory variables, whereas some models introduce memory implicitly through delays [31]. Model (3.4) can also be rewritten in the delay form, because the solution of the SDE system (3.4) is given by the variation-of-constants formula [45, Section 3.3]
@@
where ${\mathbf{y}}_{i,j}(0)$ is a prescribed initial condition at time $t=0$. The SDE form of the memory model (3.4) can, in general, be used to describe the ‘chemical memory’ of unicellular organisms, where internal variables ${\mathbf{y}}_{i}^{k}$ can be identified with concentrations of key chemical species of signal transduction networks [25, 26]. The behaviour of such memory models will depend on the properties of matrix $A^{(K)}\in\mathbb{R}^{K\times K}$, on its eigenvalues and eigenvectors. In this paper, we use the specific form of the general model (3.4) given by (3.5). In particular, choosing ${\mathbf{y}}_{i,j}(0)=0$ in equation (3.6), we have for $y_{i,j}^{1}\equiv y_{i,j}^{1}(t)$
@@
To connect the mathematical formulation with a biologically meaningful notion of memory duration, we characterize the effective memory length as the mean past time interval that contributes to the current motion of an individual. This quantity corresponds to the mean of the normalized version of the kernel $\kappa=\kappa(t)$,
@@
The quantity $K/\alpha$ therefore provides a direct measure of the biological memory timescale. It represents the average time period between when local crowding is perceived and when it influences the agent’s movement. Increasing the number of memory layers $K$ proportionally increases this period, meaning that individuals respond to a longer history of past environmental conditions rather than primarily to the recent past. From a biological perspective, larger values of $K$ correspond to organisms with slower adaptation or longer-lasting internal states, while smaller $K$ corresponds to short-term memory dominated by recent sensory input.
In Sections 6 and 7, we will numerically investigate the impact of the choice of $K$ on the clustering properties of the individual-based model with memory given by equations (2.1) and (3.9). Let us note that the kernel (3.8) takes very different shapes even if the length (3.10) is kept constant. For instance, in Figure 1 we plot $\kappa=\kappa(t)$ for the length
@@
Finally, let us note that this individual-based model can be also viewed for $K=1$ as a simplified version of what is called ‘the second-order model’ in reference [8]. Indeed, if $K=1$, then the SDE system (3.1) reduces to
@@ 4 Formal derivation of the Fokker-Planck equation
The formal large population limit as $N\to\infty$ of the system (3.1) with the average density given by (2.1) can be derived by a generalization of the procedure used in [8, Section 3]. Since some nontrivial steps were omitted in [8], we include the detailed formal derivation here. The main idea is to ‘linearize’ (3.1) by introducing the average densities $\vartheta_{i}=\vartheta_{i}(t)$ for $i\in[N]$ as a set of $N$ independent variables. According to (2.1) and the first equation of (3.1), these new variables are subject to the system of differential equations
@@
One then applies the Itô formula [45] to write down the Liouville equation [13] for the $N$-particle density $f^{N}=f^{N}(t,{\mathbf{x}}_{1},\dots,{\mathbf{x}}_{N},{\mathbf{y}}^{1}_{1},\dots,{\mathbf{y}}^{1}_{N},\dots,{\mathbf{y}}^{K}_{1},\dots,{\mathbf{y}}^{K}_{N},\vartheta_{1},\dots,\vartheta_{N})$,